# R1-Code-Interpreter: LLMs Reason with Code via Supervised and Multi-stage Reinforcement Learning

### v2

@@ C.8 Conclusions
  • Why potential-guided curriculum works. Selecting by $\Pi_{i}=4\hat{p}_{i}(1-\hat{p}_{i})$ targets the high-variance region where learning signal is largest (Prop. 2) and maintains it across stages, improving the expected per-step gain (Eq. equation C.5).
Appendix D Ablation study
Figure 9: Comparison of SFT training across different frameworks: multi-turn Code Interpreter (CI), single-turn All Text, and single-turn All Code. All frameworks are trained on the same amount of 6.5k sampled correct trajectories from GPT-4o, covering 107 training tasks. Prior to training, CI performs comparably to the other two frameworks, but after training it clearly outperforms them on both training and testing tasks. In testing tasks, SFT training on All Text and All Code yields only marginal improvements, indicating that the model overfits to the training tasks. These results demonstrate that the multi-turn Code Interpreter framework is more generalizable and effective.
@@ Appendix D Ablation study
Appendix E Description of reasoning and planning tasks
@@
T116 - Polynomial Equations The task is to find real values of a variable in a polynimial equation.
@@
Reasoning
Logical Deduction

### v3

@@ D.8 Conclusions
  • Why potential-guided curriculum works. Selecting by $\Pi_{i}=4\hat{p}_{i}(1-\hat{p}_{i})$ targets the high-variance region where learning signal is largest (Prop. 2) and maintains it across stages, improving the expected per-step gain (Eq. equation D.5).
Appendix E Performance of R1-Code-Interpreter in out-of-distribution (OOD) tasks
Table 4: Performance of R1-CI-14B and R1-CI-7B in OOD tasks: Graduate-Level Google-Proof Q&A (GPQA, Diamond) (Rein et al., 2024), and American Invitational Mathematics Examination (AIME 24&25).
Each slot is the average of three repeated runs
Model Qwen-2.5-7B Qwen-2.5-7B R1-CI-7B Qwen-2.5-14B Qwen-2.5-14B R1-CI-14B
All Text CodeSteer All Text CodeSteer
GPQA 31.2 32.9 39.0 40.1 41.2 50.2
AIME 2024 & 2025 8.33 8.33 15.0 30.0 33.3 42.0
Figure 13: Evaluation on out-of-distribution (OOD) tasks. We trained the 14B model using two-stage GRPO with SymBench and Reasoning-Gym as training data, and evaluated it on Big-Bench-Hard(BBH) as an OOD benchmark. The number of training tasks remained 107. The blue curve shows the steadily increasing reward during two-stage GRPO training. Comparing the BBH test results of the new model (green curve) with the original R1-CI-14B model (purple curve) shows similar performance, indicating that our training framework generalizes effectively to unseen task distributions from varied sources.
Appendix F Impact of sample improvement potential on GRPO training
Figure 14: GRPO training using datasets grouped by improvement potential: Group 1 [0.64–1.00], Group 2 [0.48–0.64], and Group 3 [0.32–0.48]. Each group contains the same number of samples drawn from the same Reasoning Gym and SymBench tasks, but with different improvement potential ranges. Models were evaluated on BBH for fair comparison. Models trained on higher-potential samples show consistently rising training rewards and achieve higher BBH test scores.
Appendix G Ablation studies on warm start vs. cold start, Qwen vs. DeepSeek-distilled reasoning models as base models
@@ Appendix F Impact of sample improvement potential on GRPO training
Appendix H Description of reasoning and planning tasks
@@
T116 - Polynomial Equations The task is to find real values of a variable in a polynomial equation.
@@
Reasoning Logical Deduction